OpenStudy (anonymous):

sole this limit :

OpenStudy (anonymous):

\[\lim_{x \rightarrow -2} \frac{ x^3+2x^2-4x-8 }{ x+2 }\]

OpenStudy (anonymous):

direct subst. 0/0

hartnn (hartnn):

factor the numerator

OpenStudy (anonymous):

(x+2)(x^2+4x+4)

hartnn (hartnn):

nopes

OpenStudy (anonymous):

ohh then i don't know :S

hartnn (hartnn):

take x^2 common from 1st 2 terms

OpenStudy (anonymous):

ohh, are we grouping like the last question ?

hartnn (hartnn):

yes !

hartnn (hartnn):

so ?

OpenStudy (anonymous):

would the final answer equal zero ?

hartnn (hartnn):

yup. did u understand how ?

OpenStudy (anonymous):

yupp, same as last one thanks again :)

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

but i have a question: what would happen if the trinomial didn't group what would i do ?

hartnn (hartnn):

remember this : IF YOU GET 0 AFTER DIRECT SUBSTITUTION, THEN YOU MUST BE ABLE TO FACTOR IT OUT.

OpenStudy (anonymous):

\[\lim_{x \rightarrow 2} \frac{ 2x^3-5x^2+3x-2 }{ 2x-4 }\]

hartnn (hartnn):

if for lim x->a you get a polynomial in numerator =0 when directly substituted, implies x-a MUST be factor of that polynomial

OpenStudy (anonymous):

^ i get 0/0 but in order to factor out is it okay if i need to use \[p \in { ... }\]

hartnn (hartnn):

p ?

OpenStudy (anonymous):

values of the last variable

hartnn (hartnn):

sorry, still i can't understand what p is ? for the above problem, what is p ?

OpenStudy (anonymous):

it would be all the factors of 2

OpenStudy (anonymous):

that when substituted would make it zero

hartnn (hartnn):

x-2 must be factor of 2x^3-5x^2+3x-2 because u get 0 after u put x=2

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

then do i use synthetic division ?

hartnn (hartnn):

yup.

hartnn (hartnn):

u get a quadratic, where u directly put x=2

OpenStudy (anonymous):

\[\lim_{x \rightarrow 2}\frac{ (2x^2-x-1)(x-2) }{ 2x-2 }\]

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

did u miss out 2 or 4 ?

OpenStudy (anonymous):

yeah the denominator is 2x-4

hartnn (hartnn):

i think there is an error in synthetic division, check it

OpenStudy (anonymous):

\[\lim_{x \rightarrow 2}\frac{ 2x^2-x+1)(x-2) }{ 2x-4 }\]

hartnn (hartnn):

yup, now u can cancel x-2 from numerator and denominator

OpenStudy (anonymous):

\[\frac{ 2x^2-x+1 }{ x+2 }\]

hartnn (hartnn):

what ? how u got x+2 in denominator ?

OpenStudy (anonymous):

(x-2)

hartnn (hartnn):

but x-2 cancels out

OpenStudy (anonymous):

yeah then im left with x+2

hartnn (hartnn):

how x+2 ?

OpenStudy (anonymous):

(x-2)/2x-4 is x-2 i think :S

hartnn (hartnn):

2x-4 = 2 (x-4) (x-2)/2 (x-2) = 1/ 2

OpenStudy (anonymous):

ohhhh :$haha woops

hartnn (hartnn):

put x=2 in 2x^2-x+1 / 2